OpenStudy (anonymous):

On July 16, 1882, a massive thunderstorm over Dubuque, Iowa, produced huge hailstones. The diameter of some of the hailstones was 17 in. Ice weighs about 0.033 lb/in3. What is the approximate weight of a 17 in. hailstone @Directrix

Directrix (directrix):

V = 4/3 * π * r³ Diameter = 17 so radius = 17/2 = 8.5

Directrix (directrix):

V = 4/3 * π * (8.5)³

OpenStudy (anonymous):

think i did it wrong

Directrix (directrix):

If the volume is 2572.44 cubic inches, then to get the weight is the next step.

OpenStudy (anonymous):

alright

Directrix (directrix):

2372.44 * 0.033 =

OpenStudy (anonymous):

78.29

Directrix (directrix):

Yes, to the nearest tenth that is 78.3 pounds Check your options.

OpenStudy (anonymous):

75,80,85,90

OpenStudy (anonymous):

On July 16, 1882, a massive thunderstorm over Dubuque, Iowa, produced huge hailstones. The diameter of some of the hailstones was 17 in. Ice weighs about 0.033 lb/in3. What is the approximate weight of a 17 in. hailstone? a. 75 lb c. 85 lb b. 80 lb d. 90 lb

OpenStudy (anonymous):

did i miss a piece in the original sentence ?

Directrix (directrix):

2572.44078 times 0.033 =

Directrix (directrix):

84.89 now that I allowed for more digits of pi

Directrix (directrix):

That does not seem right.

OpenStudy (anonymous):

it was 85

Directrix (directrix):

I made the error here: 2372.44 * 0.033 = I put 23 rather than 25. Sorry.

Directrix (directrix):

So, 85

OpenStudy (anonymous):

its cool .feel like im talkin to a genius or something ,like a computer

OpenStudy (anonymous):

an yes 85

Directrix (directrix):

How many more problems remain?

OpenStudy (anonymous):

well i have 10 and after that i think its a 25 question test and after that ill have to wait till 5:30 pm and take my final test to get my diploma

Directrix (directrix):

I mean how many more problems do you want to do on OpenStudy now?

OpenStudy (anonymous):

what do you mean by open study like that i can put on here ?